OpenStudy (haleyelizabeth2017):

Express 3sqrt{3} +3i in polar form.

OpenStudy (haleyelizabeth2017):

\[3\sqrt{3}+3i\]

OpenStudy (anonymous):

Hint: x+yj=r(cos θ+j sin θ) Where r is the absolute value (or modulus) of the complex number

OpenStudy (haleyelizabeth2017):

so \[6(cos\frac{\pi}{6}-i~sin\frac{\pi}{6})\]would this be correct? or...?